OpenStudy (anonymous):

How many solutions does the equation have? A. one solution B. infinite solutions C. no solution

OpenStudy (anonymous):

what equation?

OpenStudy (anonymous):

4(2x+3)=2(3x-4)

OpenStudy (anonymous):

a

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

How many solutions does the equation have? A. one solution B. infinite solutions C. no solution

OpenStudy (anonymous):

\[a +5+=\frac{ 1 }{ 5 }\left( 15 \right)\]